• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Cornejo, JC

    Log entry time 16:11:14 on January21,2008

    Entry number 226617

    keyword=Day shift summary

    On Shift:
    Juan Carlos Cornejo (SL)
    Diana Parno (TO)
    Jose Manuel (visiting)

    Shift Runplan
    1) Finish with PREX Pb target until we reach 2M events.
    2) Move to 12C target. Turn raster off and perform a bull's eye scan to test raster parameter in analysis. (per Guido's request)
    2) 12C, raster off, 1 hour @ 80uA
    3) 12C, raster on (6x4)mm^2, 1hour @ 80uA
    4) Move to thin Pb slanted target, raster on (6x4)mm^2. Increment current in 10uA/10 minutes steps until we reach 60uA. Stay at 60uA for 5 hours. This will run till next shift.


    Shift Summary:
    07:00 Running on PREX Pb target since 07:19
    08:30 Stopped run 3092. Moving to 12C target, asked MCC
    to mask target. Will start Bull's eye scan.
    08:39 Called MCC asked for 10uA beam raster off. Asked for beam
    position (0,1) center.
    08:40 Start run 3093
    08:45 Stopped run 3093 because prescales were off, producing too
    few events. Changed to 10
    08:46 Started 3094
    08:49 Stopped run 3094 because an end or run file was already
    produced while run was running.
    08:49 Started 3095
    08:56 Called MCC moved beam to (-3,1)
    09:00 Started run 3096
    09:05 Stopped 3096. Called MCC moved beam to (-5,1)
    09:13 Started run 3097
    09:19 Stopped run 3097 asked MCC to move beam to (3,1)
    09:25 Started 3098.
    09:31 Stopped run 3098 asked MCC to move beam to (5,1)
    09:35 Started run 3099.
    09:35 Stopped run 3099 hitting frame?
    09:
    10:26 Ended with Bull's eye scan. Asked MCC for 60uA beam centered
    at (0,1) and raster remains off.
    10:29 Started production run 3104 on 12C, raster off with 60uA beam.
    11:25 Called MCC asked for raster on. Their settings 9.5x5(ours 6x4)

    12:18 Stopped run 3105. Finished with carbon. Alexandre wanted to
    make an access to fix the Right spectrometer's aerogel.
    Called MCC and asked for an access. Moving target to
    thin Pb.
    13:15 Alexandre is finished asked MCC to bring back beam. Will
    start production runs on thin Pb target.

    14:15 We are now at 60uA will continue collecting Pb data till the
    end of shift

    Run Summary:
    Run# Target Current Start End Events Comments
    ---- ------ ------- ----- --- ------ --------
    3092 PREX Pb 60uA 07:19 08:30
    3093 12C 10uA 08:40 08:45 Bull's eye(0,1)
    needs change in
    prescales
    3094 12C 10uA 08:46 08:49 Bull's eye(0,1) (junk)
    3095 12C 10uA 08:49 08:56 30k Bull's eye(0,1)
    3096 12C 10uA 09:00 09:05 30k Bull's eye(-3,1)
    3097 12C 10uA 09:13 09:19 30k Bull's eye(-5,1)
    3098 12C 10uA 09:25 09:31 31k Bull's eye(3,1)
    3099 12C 10uA 09:35 09: Bull's eye(5,1)
    3100 12C 10uA 09:44 Bull's eye(0,-2)
    3101 12C 10uA Bull's eye(0,-.5)
    3102 12C 10uA 10:11 Bull's eye(0.2,5)
    3103 12C 10uA Bull's eye(0,4)

    3104 12C 60uA 10:29 11:22 1.4M
    3105 12C 60uA 11:25 12:18 1.5M

    3106 Thin Pb 10uA 13:19 Cosmics
    3107 Cosmics
    3108 Cosmics
    3109 junk
    3110 junk
    3111 Thin Pb 10uA 13:21 13:29 570k
    3112 Thin Pb 20uA 13:30 13:

    3113 - > 3121 we were ramping up current slowly and adjusting
    prescales.

    3122 Thin Pb 60uA 14:14 15:02 1M
    3123 Thin Pb 60uA 15:02 15:43 700k
    3124 Thin Pb 60uA 15:43